13) Express the complex number 2012 - Bi zer(cos(()+isiolo). Give o degrees Coos C360) in polar form:
Answer #1

e-2 - vzi Let pcoro = √2 psino = -√2 2 .82 2+224 a 822 in coro = √2 = ces 14 ces (21/4) = cos the rino - t = sin(-1/4) rin (2